KjSaif (kjsaif):

Multiply (4.7 x 10^4) by (6.4 x 10^15). Express your answer in scientific notation. A) 3.008 x 10^20 B) 3.008 x 10^60 C) 3.008 x 10^61 D) 30.08 x 10^19

OpenStudy (jiteshmeghwal9):

\[(4.7 \times 10^{-4})(6.4 \times 10^{-15})\]\[=(4.7 \times 6.4) \times (10^{-15} \times 10^{-4})\]

OpenStudy (jiteshmeghwal9):

can u solve it now ?

KjSaif (kjsaif):

yes thank you. :)

OpenStudy (jiteshmeghwal9):

The exponent in scientific notation is equal to the number of times the decimal point must be moved to produce a number between 1 and 10.

OpenStudy (jiteshmeghwal9):

yw :)